### v1

@@ 2.1 NeRF Review
Figure 2: Overview of NeRF. Figure taken from [1]
@@ 4.4 Datasets and Evaluations
For object-level datasets, ToyBox5 and ToyBox13 are two datasets proposed in NeSF [106]. Those datasets are designed to imitate scenes of children’s bedrooms and are challenging. Scenes are constructed from a large vocabulary of ShapeNet [125] objects. Each scene consists of 4-12 randomly-placed objects. Frames are rendered from 300 independently sampled camera poses. Objects are sampled at random from the 5 and 13 most common object categories, respectively for ToyBox5 and ToyBox13. PartNet [126] dataset and ShapeNet [125] are involved in evaluating the object-level segmentation performance of NeRF-based models. PartNet dataset contains a collection of object meshes that are divided into parts, along with their corresponding part segmentation labels. However, the PartNet dataset lacks the RGB information for the mesh, being not suitable for NeRF-based settings. To overcome this issue, SegNeRF [105] finds correspondences between the PartNet dataset and the ShapeNet dataset and extracts the color information for those matched object meshes from the ShapeNet dataset.
@@ 6.1 RGB-D SLAM
iMAP [183] emerged as a pioneering work, merging RGB-D SLAM with NeRF to enable real-time 3D mapping and camera tracking in an end-to-end manner. iMAP adeptly captures and reconstructs evolving scenes by incrementally optimizing a NeRF model to estimate density and color as geometry and appearance representation. After optimization of the geometry and appearance, it inversely optimizes the input pose of the NeRF model for updating the tracking results. However, iMAP’s reliance on a single MLP for neural mapping introduces memory issues as the mapping scale increases, leading to a forgetting problem. Therefore, NICE-SLAM [8], inspired by iMAP, enhences the original approach by utilizing voxel grid features to model the scene, decoding them into occupancies with pre-trained MLPs. However, the generalizability of NICE-SLAM to novel scenes is limited because of the frozen pre-trained MLPs. Another issue is the cubic memory growth rate of their model, which results in using low-resolution voxel grids and losing fine geometry details.
@@ 6.2 Monocular SLAM
Also drawing inspiration from NICE-SLAM, Li et al. [210] propose an approach that uses a multi-resolution feature grid to represent the scene. In contrast to NICER-SLAM, which utilizes an off-the-shelf monocular predictor for depth prediction and enforces depth consistency between the rendered expected depths and monocular depths, this method incorporates a multi-view stereo loss to enforce consistent depth and color predictions among closely positioned images to jointly update camera poses and the map. However, a potential concern arises as the MLP decoder is fixed after the initialization stage, which might cause problems when the camera moves to a region significantly different from those in the initialization.
@@
For example, VIO NeRF [211] leverage motion priors obtained from Apple ARKit as initial camera poses to fit the NeRF network and then jointly optimize 3D scene representations and camera poses.
@@
NeRF-VO [216] chooses a learning-based sparse visual odometry [217] for camera tracking and optimizes a NeRF scene representation and camera poses by incorporating dense depth and surface normal supervision. NeRF-SLAM [218] also utilizes a learning-based visual odometry, Droid-SLAM [219] to estimate initial accurate poses and depth maps with associated uncertainty. The objective of NeRF-SLAM is to achieve a real-time estimation of both geometric and photometric maps of the scene by leveraging techniques such as probabilistic volumetric fusion and InstantNGP (hash-based hierarchical volumetric radiance fields) [220]. However, NeRF-SLAM only performs offline global bundle adjustment at the end of camera tracking. In certain challenging scenarios, it’s difficult to eliminate drift errors solely through offline refinement, as it lacks online loop closure detection and global bundle adjustment. This limitation restricts its ability to achieve globally consistent 3D reconstruction. Therefore, in addition to employing Droid-SLAM as the frontend for tracking, HI-SLAM [221] introduces an SDF to enhance surface definition and incorporates loop closing to ensure global consistency and GO-SLAM [222] introduce an efficient loop closing and online full bundle adjustment.
@@ 6.3 LiDAR SLAM
LONER [223] first introduces the use of a neural implicit scene representation in the LiDAR SLAM system. The tracking module processes incoming LiDAR scans and estimates odometry using point-to-plane ICP. The scene is represented as an MLP with hierarchical feature grid encoding. To ensure depth supervision and prevent the NeRF model from forgetting learned geometry and experiencing slow convergence when integrating new LiDAR scans, the author optimizes the MLP using a dynamic margin loss combined with depth and sky losses (where weights are set to zero for rays pointing at the sky). EINRUL [224] does not utilize depth supervision due to its investigation into the discrepancy between LiDAR frames and depth images. It finds that a LiDAR frame contains only 5% of the depth data found in a depth image, indicating that depth rendering may not offer sufficient geometry information for supervision. Therefore, EINRUL opts to supervise directly in 3D space, eliminating the need for volume rendering and ground truth labels. This approach allows for the refinement of initial poses while optimizing the implicit representation.
@@
Since LiDAR SLAM is mainly utilized in outdoor environments, whereas NeRFs often struggle with outdoor, unbounded scenes. Addressing incremental LiDAR inputs with NeRF within large-scale environments presents a challenge. In response, an octree is commonly used to recursively divide the scene into leaf nodes, which consist of basic scene units known as voxels, enabling incremental representation of large-scale outdoor environments. As a result, NeRF-LOAM [225] utilizes octree-based voxels in combination with neural implicit embeddings. These embeddings are decoded into a continuous SDF by a neural implicit decoder. The embeddings, decoder, and poses are optimized concurrently by minimizing SDF errors. NF-Atlas [226] adopts the concept of dividing the large-scale scene into multiple submaps using multi-volume neural feature fields. In this approach, neural feature volumes serve as pose graph nodes, with the relative pose between volumes acting as pose graph edges, ensuring both local rigidity and global elasticity of the entire neural feature field. Locally, NF-Atlas employs a sparse feature octree and a compact MLP to encode the SDF of these submaps, facilitating end-to-end Maximum a Posteriori (MAP) based probabilistic mapping. Transitioning to CLONeR [227], this system introduces occupancy grids to efficiently bypass ray marching in areas devoid of objects within unbounded scenes. CLONeR utilizes two separate MLPs: one to learn occupancy from LiDAR ray data and another to model the scene’s RGB characteristics from camera rays. This dual-MLP configuration helps decouple the occupancy and visual data processing, ensuring that visual artifacts or illumination variances in the RGB images do not compromise the accuracy of the geometric information, such as the depth rendered from LiDAR data.
@@
Maintaining global map consistency is also vital for improving reconstruction accuracy in large-scale environments. In this regard, NF-Atlas adjusts the origins of its volumes to align with the revised robot trajectory when a loop closure occurs. On the other hand, PIN-SLAM [228] adopts a different approach, utilizing a Point-Based Implicit Neural Representation to achieve global map consistency.
@@
Reconstructing high-accuracy maps with resource constrained devices has also emerged as a key area of research. Shi et al. [229] tackle the challenge of generating precise maps in memory-limited hardware environments. They initiate the process by allocating a spatial hash grid near geometric surfaces using aligned batch range data. Points along rays within the grid are then sampled simultaneously. Each sampled point is encoded into a feature vector using a hierarchical hash encoder and subsequently transformed into a TSDF value by a small MLP.
@@ 6.4 Datasets and Evaluations
Datasets: When evaluating the above-mentioned methods, a large number of datasets are used. Regarding NeRF-based visual SLAM methods, three principle datasets stand out because of their extensive use in this field: Replica[86], TUM[230], and ScanNet[87]. These three datasets are particularly favored for providing comprehensive RGB and Depth images, as well as the ground truth poses. More importantly, these datasets focus on the indoor environment, which not only facilitates a detailed evaluation of mapping and tracking accuracy but also serves as a perfect scenario for testing the ability of novel view synthesis. In particular, Replica and ScanNet datasets have been introduced in Section 3.3, while the TUM dataset includes a variety of indoor scenes such as offices, a technical lab, household environments, and a large lecture room. For each sequence, the dataset provides synchronized RGB and depth images, camera intrinsics, and ground truth poses. The ground truth data is obtained using a high-accuracy motion-capture system in smaller rooms or a high-precision laser tracker in larger rooms.
@@ 7.1 Navigation
Visual localization and navigation are important tasks in robotics — given a query image, finding the location of this image in the environment, and navigating to reach the pose that has the same view as the query image. RNR-Map [239] addresses those challenges by leveraging a latent-code-based NeRF [240]. Similar to conventional occupancy maps, the constructed map has a grid form but the information stored in each grid is not the occupancy but a latent code, which is used as an intermediate representation of the image at the corresponding place. The latent code is converted from the input image by an encoder and is decoded to an input to the NeRF model [240] by a decoder. Different from mobile navigation in indoor environments [239], the task of visual localization and navigation can also happen to manipulators, which are considered in NSR-VC, where a time contrastive loss is combined with NeRF in an autoencoding framework to learn viewpoint-invariant 3D-aware scene representations, which further enable visuomotor control for challenging manipulation tasks. Similar to the latent-code-based NeRF [239, 240], Comp-NeRF uses latent vectors to parameterize individual NeRFs from which the scene can be reconstructed. Then a graph neural network is trained in the latent space to capture compositionality in dynamic multi-object scenes.
@@ 7.3 Datasets and Evaluations
Datasets: Several public image-goal navigation datasets are involved. NRNS [252] dataset contains three navigation difficulty levels (easy, medium, hard) with two path categories (straight and curved). Each level has 1000 episodes for each path type, except for the hard-straight set (only 806). Gibson [253] dataset has 72 houses for the training set and 14 houses for the validation set. MP3D [254] is a comprehensive RGB-D dataset comprising 10,800 panoramic views from 194,400 RGB-D images across 90 building-scale scenes. It includes annotations for surface reconstructions, camera poses, and both 2D and 3D semantic segmentations. Among all the reviewed papers, some works, such as NSR-VC [255], leverage existing simulators, e.g., NVIDIA FleX [256] to generate environments for experiments. Other simulators used include PyBullet [257] and Habitat [258].

### v2

@@ 2.1 NeRF Review
Figure 2: Overview of NeRF. Figure modified from [1]
@@ 4.4 Datasets and Evaluations
For object-level datasets, ToyBox5 and ToyBox13 are two challenging datasets proposed in NeSF [106], which mimic scenes of children’s bedrooms using objects from ShapeNet [125]. There are 4-12 randomly-selected objects randomly-placed in each scene. Additionally, 300 independently sampled camera poses are used to render input frames. PartNet [126] dataset and ShapeNet [125] are involved in evaluating the object-level segmentation performance of NeRF-based models. PartNet dataset contains a collection of object meshes that are divided into parts, along with their corresponding part segmentation labels. However, the PartNet dataset lacks the RGB information for the mesh, being not suitable for NeRF-based settings. To overcome this issue, SegNeRF [105] finds correspondences between the PartNet dataset and the ShapeNet dataset and extracts the color information for those matched object meshes from the ShapeNet dataset.
@@ 6.1 RGB-D SLAM
iMAP [183] emerged as a pioneering work, merging RGB-D SLAM with NeRF to enable real-time 3D mapping and camera tracking in an end-to-end manner. iMAP adeptly captures and reconstructs evolving scenes by incrementally optimizing a NeRF model to estimate density and color as geometry and appearance representation. After optimization of the geometry and appearance, it inversely optimizes the input pose of the NeRF model for updating the tracking results. However, iMAP’s reliance on a single MLP for neural mapping introduces memory issues as the mapping scale increases, leading to a forgetting problem. Therefore, NICE-SLAM [8], inspired by iMAP, improves the original method by employing voxel grid features to model the scene and decoding these features into occupancies using pre-trained MLPs. However, NICE-SLAM’s ability to generalize to novel scenes is constrained due to the fixed pre-trained MLPs. In addition, the model’s cubic memory growth rate tends to use low-resolution voxel grids, resulting in the loss of fine geometric details.
@@ 6.2 Monocular SLAM
Also drawing inspiration from NICE-SLAM, Dense RGB SLAM [210] propose an approach that uses a multi-resolution feature grid to represent the scene. In contrast to NICER-SLAM, which utilizes an off-the-shelf monocular predictor for depth prediction and enforces depth consistency between the rendered expected depths and monocular depths, this method incorporates a multi-view stereo loss to enforce consistent depth and color predictions among closely positioned images to jointly update camera poses and the map. However, a potential concern arises as the MLP decoder is fixed after the initialization stage, which might cause problems when the camera moves to a region significantly different from those in the initialization.
@@
For example, VIO NeRF [211] utilizes Apple ARKit to obtain the motion priors as initial camera poses to fit the NeRF network and then jointly optimize 3D scene representations and camera poses.
@@
NeRF-VO [216] adopts a learning-based method in sparse visual odometry [217] to estimate pose tracking and incorporate dense depth and surface normal supervision in a joint optimization between NeRF scene representation and camera poses. NeRF-SLAM [218] also utilizes a learning-based visual odometry, Droid-SLAM [219] to estimate initial accurate poses and depth maps with associated uncertainty. The objective of NeRF-SLAM is to achieve a real-time estimation of both geometric and photometric maps of the scene by leveraging techniques such as probabilistic volumetric fusion and InstantNGP (hash-based hierarchical volumetric radiance fields) [220]. However, NeRF-SLAM only performs offline global bundle adjustment at the end of camera tracking. In certain challenging scenarios, it’s difficult to eliminate drift errors solely through offline refinement, as it lacks online loop closure detection and global bundle adjustment. This limitation restricts its ability to achieve globally consistent 3D reconstruction. Therefore, in addition to employing Droid-SLAM as the frontend for tracking, HI-SLAM [221] introduces an SDF to enhance surface definition and incorporates loop closing to ensure global consistency and GO-SLAM [222] introduce an efficient loop closing and online full bundle adjustment.
@@ 6.3 LiDAR SLAM
LONER [223] first introduces the use of a neural implicit scene representation in the LiDAR SLAM system. The tracking module processes incoming LiDAR scans and estimates odometry using point-to-plane ICP. The scene is represented as an MLP with hierarchical feature grid encoding. To ensure depth supervision and prevent the NeRF model from forgetting learned geometry and experiencing slow convergence when integrating new LiDAR scans, the author optimizes the MLP using a dynamic margin loss combined with depth and sky losses (where weights are set to zero for rays pointing at the sky). EINRUL [224] does not utilize depth supervision due to its investigation into the discrepancy between LiDAR frames and depth images. It finds that a LiDAR frame contains only 5% of the depth data found in a depth image, indicating that depth rendering may not offer sufficient geometry information for supervision. Therefore, EINRUL opts to supervise 3D space directly, eliminating the need for volume rendering and ground truth labels. This approach enables the refinement of initial poses alongside the optimization of the implicit representation.
@@
Since LiDAR SLAM is mainly utilized in outdoor environments, whereas NeRFs often struggle with outdoor, unbounded scenes. Addressing incremental LiDAR inputs with NeRF within large-scale environments presents a challenge. In response, an octree is typically employed to recursively split the scene into leaf nodes, each containing basic scene units known as voxels, enabling incremental representation of large-scale outdoor environments. Consequently, NeRF-LOAM [225] employs octree-based voxels alongside with neural implicit embeddings. These embeddings are converted into a continuous SDF by a neural implicit decoder. The joint optimization among embeddings, decoders, and poses is carried out by minimizing SDF errors. NF-Atlas [226] adopts the concept of dividing the large-scale scene into multiple submaps using multi-resolution neural feature grids. In this approach, neural feature grids serve as nodes of the pose graph, while the relative pose between feature grids acts as the edges of the pose graph. This method ensures both local rigidity and global elasticity of the entire neural feature grid. Locally, NF-Atlas deploys a sparse feature octree and a compact MLP to encode the SDF of these submaps, enabling an end-to-end probabilistic mapping using Maximum a Posteriori (MAP). Transitioning to CLONeR [227], this system introduces occupancy grids to efficiently bypass ray marching in areas devoid of objects within unbounded scenes. CLONeR utilizes two separate MLPs: one to learn occupancy from LiDAR ray data and another to model the scene’s RGB characteristics from camera rays. This dual-MLP configuration helps decouple the occupancy and visual data processing, ensuring that visual artifacts or illumination variances in the RGB images do not compromise the accuracy of the geometric information, such as the depth rendered from LiDAR data.
@@
Maintaining global map consistency is also vital for improving reconstruction accuracy in large-scale environments. In this regard, NF-Atlas adjusts the volume origins to align with the robot trajectory revised by the occurrence of loop closure. On the other hand, PIN-SLAM [228] adopts a different approach, utilizing a Point-Based Implicit Neural Representation to achieve global map consistency.
@@
Reconstructing high-accuracy maps with resource-constrained devices has also emerged as a key area of research. Shi et al. [229] tackle the challenge of generating precise maps in memory-limited hardware environments. They initiate the process by allocating a spatial hash grid near geometric surfaces using aligned batch range data. Points along rays within the grid are then sampled simultaneously. Each sampled point is transformed into a feature vector using a hierarchical hash encoder and then encoded into a TSDF value by a compact MLP.
@@ 6.4 Datasets and Evaluations
Datasets: When evaluating the above-mentioned methods, a large number of datasets are used. Regarding NeRF-based visual SLAM methods, three principle datasets stand out because of their extensive use in this field: Replica[86], TUM[230], and ScanNet[87]. These three datasets are particularly favored for providing comprehensive RGB and Depth images, as well as the ground truth poses. More importantly, these datasets focus on the indoor environment, which not only facilitates a detailed evaluation of mapping and tracking accuracy but also serves as a perfect scenario for testing the ability of novel view synthesis. In particular, Replica and ScanNet datasets have been introduced in Section 3.3, while the TUM dataset includes a variety of indoor scenes such as offices, a technical lab, household environments, and a large lecture room. For each sequence, the dataset provides synchronized RGB and depth images, camera intrinsics, and ground truth poses. The ground truth data is obtained using a motion-capture system with highly accurate resolution in smaller rooms or a high-precision laser tracker in larger rooms.
@@ 7.1 Navigation
Visual localization and navigation are important tasks in robotics — given a query image, finding the location of this image in the environment, and navigating to reach the pose that has the same view as the query image. RNR-Map [239] addresses those challenges by leveraging a latent-code-based NeRF [240]. Similar to conventional occupancy maps, the constructed map has a grid form but the information stored in each grid is not the occupancy but a latent code, which is used as an intermediate representation of the image at the corresponding place. The latent code is converted from the input image by an encoder and is decoded to an input to the NeRF model [240] by a decoder. Different from mobile navigation in indoor environments [239], the task of visual localization and navigation can also happen to manipulators, as in NSR-VC. Specifically, a time contrastive loss is combined with NeRF in an autoencoding framework to learn viewpoint-invariant scene representations. Such a representation is later used to enable visuomotor control for challenging manipulation tasks. Similar to the latent-code-based NeRF [239, 240], Comp-NeRF uses latent vectors to parameterize individual NeRFs from which the scene can be reconstructed. Then a graph neural network is trained in the latent space to capture compositionality in dynamic multi-object scenes.
@@ 7.3 Datasets and Evaluations
Datasets: Several public image-goal navigation datasets are involved. NRNS [252] dataset contains straight and curved paths, which are classified into easy, medium, and hard, depending on the navigation difficulty. Apart from the hard-straight set, which only has 806 episodes, each difficulty-path type combination encloses 1000 episodes. Gibson [253] dataset, as another popular in-house dataset for navigation tasks, has 86 houses in total, with 72 for training and 14 for validation. In addition, MP3D [254] is a comprehensive RGB-D dataset. It covers 90 building-scale scenes and provides 10,800 panoramic views from 194,400 RGB-D images. It also provides annotations for various tasks, including 3D reconstructions, camera pose estimation, and 2D and 3D semantic segmentation. Among all the reviewed papers, some works, such as NSR-VC [255], leverage existing simulators, e.g., NVIDIA FleX [256] to generate environments for experiments. Other simulators used include PyBullet [257] and Habitat [258].